Find the distance between the following pair of points
A) the origin and (-√3,√6)
1
Expert's answer
2020-08-30T18:03:01-0400

The distance , d , between two points (x1,y1) and (x2,y2) is given by d = "\sqrt{(x_1-x_2)^2 +(y_1-y_2)^2}" So, here we have (x1,y1)=(0,0) and (x2,y2)=("-\sqrt3 ,\sqrt6" )Hence distance is d = "\sqrt{\smash[b](0-(-\sqrt3))^2+(0-\sqrt6)^2}" = "\sqrt{\smash[b](\sqrt3)^2+(\sqrt6)^2}"
=> d = "\sqrt{\smash[b]3+6} = \sqrt{\smash[b]9} = 3" units.